Use the addition identity to find the exact value for sin 105°. Show the work.
Answer by lwsshak3(11628)

Use the addition identity to find the exact value for sin 105°. Show the work.
..
Addition formula for sin: sin(s+t)=sin s cos t+cos s sin t
For given problem:
sin 105º=sin(45+60)=sin 45 cos60+ cos 45 sin 60
=√2/2*1/2+√2/2*√3/2
=√2/4+√6/4
=(√2+√6)/4 (ans)